OpenStudy (anonymous):

what is the half of 2 4/5, 5/6 and 3 1/2? please help

OpenStudy (anonymous):

half of 2 4/5 = 1/2 * 2 4/5 2 4/5 = (5*2) + 4 / 5 = 14/5 1/2 * 14/5 = 7/5 Do all the problems same way

OpenStudy (anonymous):

half of 2 4/5 is 7/5, half of 5/6 is 5/12, and half of 3 1/2 is 1 3/4.